# GLOV: Guided Large Language Models as Implicit Optimizers for Vision Language Models

### v1

@@ Abstract
In this work, we propose a novel method (GLOV) enabling Large Language Models (LLMs) to act as implicit Optimizers for Vision-Language Models (VLMs) to enhance downstream vision tasks. Our GLOV meta-prompts an LLM with the downstream task description, querying it for suitable VLM prompts (e.g., for zero-shot classification with CLIP). These prompts are ranked according to their fitness for the downstream vision task. In each respective optimization step, the ranked prompts are fed as in-context examples (with their accuracies) to equip the LLM with the knowledge of the type of prompts preferred by the downstream VLM. Furthermore, we also explicitly steer the LLM generation in each optimization step by specifically adding an offset difference vector of the embeddings from the positive and negative solutions found by the LLM, in previous optimization steps, to the intermediate layer of the network for the next generation step. This offset vector steers the LLM generation toward the type of language preferred by the downstream VLM, resulting in enhanced performance on the downstream vision tasks. We comprehensively evaluate our GLOV on 16 diverse datasets using two families of VLMs, i.e., dual-encoder (e.g., CLIP) and encoder-decoder (e.g., LLaVa) models – showing that the discovered solutions can enhance the recognition performance by up to $15.0\%$ and $57.5\%$ ($3.8\%$ and $21.6\%$ on average) for these models.
00footnotetext: Work partially done as an intern at Sony Group Corporation, Japan.
$0$.Correspondence: mirza@tugraz.at
$0$.
Figure 1: The effect of prompt evolution on the downstream task performance. The shaded regions represent the absolute top-1 classification accuracies for ImageNet (Deng et al., 2009) at each optimization step by ensembling the top-3 prompts found w.r.t the accuracy on the 1-shot train set whereas the solid lines represent the exponential moving average. The left plot is with CLIP VIT-B/32 (Radford et al., 2021), and the right is with LLaVa-OV (Li et al., 2024) while the LLM employed is Llama-3 (Dubey et al., 2024). Due to high computational cost, we only perform $25$ optimization steps for LLaVa-OV.
@@ 1 Introduction
Orthogonal to traditional gradient-based optimization (Nesterov, 1983, Boyd & Vandenberghe, 2004, Kingma & Ba, 2014, Ruder, 2016), the recent rise of large language models (Brown et al., 2020, OpenAI, 2023, Chiang et al., 2023, Raffel et al., 2020, Touvron et al., 2023a; b, Dubey et al., 2024) and vision-language foundation models (OpenAI, 2023, Li et al., 2024, Zhu et al., 2024, Alayrac et al., 2022, Radford et al., 2021) has introduced the possibility of framing optimization in the context of natural language prompts. This form of optimization typically does not require any gradient-based learning or parameter update but focuses on extracting knowledge from the language models via suitable natural language prompts. A large body of work focuses on finding natural language prompts optimized for various downstream tasks for both LLMs (Yang et al., 2024, Wei et al., 2022, Kojima et al., 2022, Yao et al., 2023) and VLMs (Pratt et al., 2023, Roth et al., 2023, Mirza et al., 2024), demonstrating impressive gains in language-based and downstream vision tasks.
@@
In our work, we frame optimization around discovering suitable natural language prompts for VLMs, with the objective of improving performance on downstream vision tasks. Our proposed GLOV employs a prompt search technique relying on a meta-prompt coupled with embedding space guidance, that drives the prompt optimization for the VLMs. We use the meta-prompt to iteratively query an LLM with downstream task-specific description and ranked in-context examples derived from the previous (optimized) prompts. In-context examples guide the LLM toward the desired output, and their ranking (measured on a small held-out train set) provides the LLM with a sense of the language patterns preferred by the downstream VLM. To further steer the LLM generation towards a notion of goodness in each optimization step, we explicitly bias the language generation with a direction. The direction is determined by adding a hidden state offset vector (on the last token during the autoregressive generation) derived from the positive and negative prompts (based on their effectiveness on labeled training data) to the LLM’s activation space during generation. The intuition is that by directing the LLM generation toward the positive prompts, the model can discover semantically similar and potentially more effective solutions. One complete optimization run, depicting the effectiveness of the discovered solutions and the effect of applying the embedding space guidance is plotted in Figure 1. The best-performing prompts (on the held-out train set) achieve an absolute improvement of $2.6\%$ and $15.2\%$ on ImageNet (Deng et al., 2009) test set over CLIP (Radford et al., 2021) and LLaVa-OV (Li et al., 2024) respectively.
@@
We extensively evaluate our GLOV on one of the fundamental tasks in computer vision: image classification, and also touch upon the open-ended generation task of visual question answering (VQA). We demonstrate the generalization of our GLOV on a total of $16$ diverse datasets, with the two commonly employed families of VLM models – the dual-encoder and the recent visual encoder-decoder models (Radford et al., 2021, Li et al., 2024). We find that our GLOV can consistently discover highly effective solutions for the downstream task of interest resulting in significant improvements across the board. For example, the most effective prompts discovered for the dual-encoder models (e.g., CLIP) can improve the accuracy up to $15.0\%$ ($3.81\%$ on average) and for the encoder-decoder architectures (e.g., LLaVa), the resulting prompts show an even larger improvement of up to $57.5\%$ ($21.6\%$ on average). Furthermore, we extensively ablate our proposed prompt optimization algorithm, design choices, and the effect of our guidance mechanism, providing insights for future work.
@@ 2 Related Work
Our work is related to large language and vision language models, approaches proposing methods for steering the LLM outputs, and prompt optimization methods (through LLMs) for VLMs.
@@ 2.1 LLMs and VLMs
Here, we first provide a brief overview of Large-Language Models (LLMs) and then move towards Vision-Language Models (VLMs).
@@
LLMs have revolutionized the natural language processing landscape. These models can typically be divided into two major groups; namely the long-short-term-memory (LSTM) (Hochreiter & Schmidhuber, 1997) and transformer-based architectures (Vaswani et al., 2017). The former is based upon recurrent neural networks (RNNs) that use gates to control the flow of information, allowing them to capture long-term dependencies in sequential data. The latter is based on the self-attention mechanism, which enables the model to process sequences in parallel and capture relationships between tokens regardless of distance. Some notable works following the LSTM family of models include (Sutskever et al., 2014, Graves & Schmidhuber, 2005, Bahdanau et al., 2015, Beck et al., 2024). The transformer-based architectures consist of encoder or decoder-based LLMs. The encoder-based LLMs are primarily used for understanding tasks like text classification and sentiment analysis, as they excel at capturing contextual information from the input. Some notable works include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019b), and DistilBERT (Sanh et al., 2020). The decoder-based LLMs, on the other hand, are designed for generative tasks such as text generation, translation, and summarization, with recent models including GPT-3 (Brown et al., 2020), T5 (Raffel et al., 2020), GPT-4 (OpenAI, 2023), and the Llama family of models (Dubey et al., 2024). In our work, since we need to access the weights of the LLMs, we resort to the open-source Llama-3 model. However, potentially any open-source LLM can be employed.
@@
VLMs can be placed in two categories. One group relies on dual-encoders (vision and text encoder), usually trained in a contrastive manner and these models are typically strong at tasks like image recognition. The most common among these methods are CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), OpenCLIP (Schuhmann et al., 2022), SigLIP (Zhai et al., 2023), and MetaCLIP (Xu et al., 2023). Many methods (Mirza et al., 2024; 2023b, Doveh et al., 2023b; a, Lin et al., 2023, Mirza et al., 2023a) build upon these models to further improve them for specific downstream tasks. The other group of methods aligns the visual modality with a frozen LLM and can be used for open-ended visual reasoning tasks like image captioning, visual question-answering, etc. Some representative approaches from this group include BLIP-2 (Li et al., 2023), Instruct-BLIP (Dai et al., 2023), MiniGPT (Zhu et al., 2024, Chen et al., 2024), and the LLaVa family of models (Liu et al., 2023, Li et al., 2024). Similarly, some approaches (Doveh et al., 2024, Gavrikov et al., 2024, Lin et al., 2024, Huang et al., 2024) build upon these models and provide further improvements. In our work, we focus on the task of object recognition by employing both families of models and frame the task of finding the optimal prompt templates (for CLIP (Radford et al., 2021)) and suitable prompts for open-ended generation (for LLaVa (Li et al., 2024)) as an optimization problem. Specifically, for the decoder-based VLMs, some recent works (Zhang et al., 2024) highlight that these models struggle for fine-grained object recognition. However, we show in our work, for the first time, that our GLOV can discover an optimal prompt that can greatly improve the visual recognition ability of these models, without requiring any gradient-based learning or fine-tuning.
2.2 Steering LLM Responses
One line of work alters the responses from an LLM, without requiring any explicit gradient-based fine-tuning. ActADD (Turner et al., 2023) proposes to reduce LLM hallucinations by altering the hidden states. In their work, given a positive and a negative data point (model response), they propose to steer the responses (e.g., to be less hateful) by adding the difference of the embeddings from these points to the intermediate layers of the network. On the other hand, Proxy-Tuning (Liu et al., 2024a) proposes to adapt the model responses on the logit level. Specifically, they model the responses from a smaller base model to be similar to a larger instruction-tuned LLM by altering the softmax probabilities obtained from the smaller base model. We also take inspiration from Turner et al. (2023), Liu et al. (2024a) and for the first time show that such steering (applied on the embedding level) can also be used to improve downstream multi-modal (vision-language) tasks. One difference between ActADD and our GLOV is that we add the difference of the sentence embeddings (and not of the prompts themselves, as in ActADD) to only the last token. Whereas, ActADD adds the difference of the (complete) sequence lengths to the first few tokens at the generation step. In the ablations Section 4.3, we show that this seeming nuance has an important effect on the performance of the downstream task, showing that our method might be more suitable for vision tasks.
@@ 2.3 Large-language Models as Prompt Optimizers
Some approaches propose employing LLMs (in an agentic workflow) to search for the optimal prompt for the downstream task. OPRO (Yang et al., 2024) coins the term “LLMs as Optimizers” and proposes iteratively discovering solutions (prompts) for natural language tasks by employing an LLM in a feedback loop. Similarly, Liu et al. (2024b) proposes to find suitable prompts for dual-encoder VLMs (e.g., CLIP) by iteratively prompting an LLM. Our GLOV also proposes to discover suitable prompts for VLMs but differs from Liu et al. (2024b) in the sense that we are employing a meta-prompt that captures long-range dependencies by tapping into the history-buffer of the in-context examples and employs task-specific knowledge that helps to obtain prompts better suited for the downstream task. Furthermore, we propose a new method for steering the LLM generation (through embedding space guidance) towards the responses that are more suitable for downstream VLMs. Powered by a suitable meta-prompt and the guidance scheme, our GLOV discovers solutions which help to enhance visual tasks for both the dual-encoder and encoder-decoder models.
3 GLOV: Guided LLMs as Implicit Optimizers for VLMs
The goal of our GLOV is to improve the VLM’s downstream (vision) task performance by optimizing natural language prompts through employing an LLM in an iterative workflow. To achieve this, we build upon a meta-prompt introduced by Mirza et al. (2024), differing from them, we leverage few-shot (e.g., 1-shot) held-out labeled training examples to calculate the effectiveness of the solutions discovered in each optimization step, which guides the optimization. Furthermore, effective prompt optimization is performed by providing the LLM with explicit guidance conditioned on a prior of the difference of the sentence embeddings from the positive and negative prompts discovered during the previous optimization iterations. Although the application space of GLOV is general and we demonstrate its generalization ability on two popular families of VLMs (e.g., dual-encoder (Radford et al., 2021) and encoder-decoder (Liu et al., 2023)), for simplicity, here we focus our description around CLIP (Radford et al., 2021) while mentioning the differences for LLaVa (Li et al., 2024) where appropriate. An overview of our methodology is provided in Figure 2.
For the ease of assimilation, we divide the description of our GLOV into different parts. In Section 3.1 we describe the fitness function and how it can provide an interface for the LLM-VLM interaction. In Section 3.2, we provide details about the meta-prompt employed in our work. Finally, we conclude in Section 3.3 by providing details about the proposed guidance methodology.
@@ 3.1 LLM-VLM interaction through Fitness Function
Figure 2: Overview of GLOV. GLOV consists of a Meta Prompt, which constitutes system instruction, task description, and in-context examples (VLM prompts) which are evaluated (and ranked) on a few-shot training data in each iteration. The Meta-Prompt instructs the LLM to generate several candidate solutions in each optimization iteration, conditioned on the in-context examples which are fed in conjunction with the accuracy values, highlighting their effectiveness. Furthermore, to steer the LLM generation towards the language preferred by the VLM, we add the scaled difference of the sentence embeddings (autoregressively) from the positive and negative text prompts to the intermediate layer of the LLM. This process is repeated until the stopping condition is met (e.g., maximum iterations). Note, that $H_{+}$ and $H_{-}$ refer to the sentence embeddings from the text prompts.
The dataset-specific prompt templates $\mathcal{P}$ provided by CLIP (Radford et al., 2021) have been constructed manually, requiring human effort. In this work, we frame the prompt search as an optimization problem and propose to replace the human with an LLM, employed in an iterative feedback loop. Furthermore, we explicitly guide the generation process of the LLM in each optimization step by proposing a novel guidance methodology that can assist the LLM in understanding the style of language preferred by the downstream VLM, even though the two models only interact through a fitness function. At each optimization step $i$, the LLM provides multiple (e.g., $10$) solutions to improve the downstream task performance. However, not all solutions provided by the LLM are preferred for the downstream vision task. To obtain a measure of the fitness (effectiveness) of the provided solutions to the downstream vision task, we evaluate all the candidate solutions on a held-out few-shot (1-shot) labeled training dataset $\mathcal{D}$. For CLIP (Radford et al., 2021), the zero-shot likelihood of class $\hat{c}$ for each discovered prompt $p\in\mathcal{P}$ during an optimization step can be found by
@@
$\text{Fitness}(p)=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}\left[\arg\max_{c}l_{c}(x)=y\right],$ (2)
where $\mathbbm{1}$ is an indicator function that is 1 if the predicted label matches the ground truth $y$ and 0 otherwise. For the encoder-decoder models (e.g., LLaVa), which produce open-ended outputs, we obtain the class likelihoods by obtaining a symbolic representation of the image in textual form and comparing the text embeddings from this symbolic representation with the text embeddings obtained for the individual class names, using a dedicated sentence embedding model (Reimers & Gurevych, 2019). We expand on these details in the Appendix Section A.
It is important to note that the fitness function forms a bridge between two disjoint models – the LLM and the VLM, and is responsible for their interaction. The fitness (classification accuracy) provides feedback to the LLM regarding the type of natural language sentences that are preferred by the downstream VLM. The fitness function is responsible for ranking all the prompts provided as in-context examples to the meta-prompt in each optimization iteration (c.f., Section 3.2) and also forms the basis for the application of the embedding-space guidance methodology (c.f., Section 3.3) proposed in this work to bias the LLM responses towards a notion of goodness.
3.2 Meta-Prompt
The meta-prompt (c.f., Appendix Figure 5) is responsible for driving the iterative prompt optimization. Specifically, it consists of $3$ distinct parts, which are described as follows:
System prompt is a generic set of instructions that describe the task of the LLM. It helps a user to find the optimal prompts, improving the downstream task accuracy. The system prompt remains static for the entire optimization run.

### v5

@@ Abstract
In this work, we propose GLOV, which enables Large Language Models (LLMs) to act as implicit optimizers for Vision-Language Models (VLMs) to enhance downstream vision tasks. GLOV prompts an LLM with the downstream task description, querying it for suitable VLM prompts (e.g., for zero-shot classification with CLIP). These prompts are ranked according to their fitness for the downstream vision task. In each respective optimization step, the ranked prompts are fed as in-context examples (with their accuracies) to equip the LLM with the knowledge of the type of prompts preferred by the downstream VLM. Furthermore, we explicitly guide the LLM’s generation at each optimization step by adding an offset vector – calculated from the embedding differences between previous positive and negative solutions – to the intermediate layer of the network for the next generation. This offset vector biases the LLM generation toward the type of language the downstream VLM prefers, resulting in enhanced performance on the downstream vision tasks. We comprehensively evaluate our GLOV on two tasks: object recognition and the critical task of enhancing VLM safety. Our GLOV shows performance improvement by up to $15.0\%$ and $57.5\%$ for dual-encoder (e.g., CLIP) and encoder-decoder (e.g., LLaVA) models for object recognition and reduces the attack success rate (ASR) on state-of-the-art VLMs by up to $60.7\%$.
Machine Learning, ICML
00footnotetext: Work partially done as an intern at Sony, Japan.
Figure 1: The effect of prompt evolution on the downstream task performance. The shaded regions represent the absolute top-1 classification accuracies for ImageNet (Deng et al., 2009) at each optimization step by ensembling the top-3 prompts found w.r.t the accuracy on the 1-shot train set whereas the solid lines represent the exponential moving average. LLM employed is Llama-3 (Dubey et al., 2024).
@@ 1 Introduction
Orthogonal to traditional gradient-based optimization (Nesterov, 1983; Boyd & Vandenberghe, 2004; Kingma & Ba, 2014; Ruder, 2016), the recent rise of large language models (Brown et al., 2020; OpenAI, 2023; Chiang et al., 2023; Raffel et al., 2020; Touvron et al., 2023a, b; Dubey et al., 2024) and vision-language foundation models (OpenAI, 2023; Li et al., 2024; Zhu et al., 2024; Alayrac et al., 2022; Radford et al., 2021) has introduced the possibility of framing optimization in the context of natural language prompts. This form of optimization typically does not require any gradient-based learning or parameter update but focuses on extracting knowledge from the language models via suitable natural language prompts. A large body of work focuses on finding natural language prompts optimized for various downstream tasks for both LLMs (Yang et al., 2024; Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) and VLMs (Pratt et al., 2023; Roth et al., 2023; Mirza et al., 2024), demonstrating impressive gains in language-based and downstream vision tasks.
@@
In our work, we frame optimization around discovering suitable natural language prompts for VLMs, with the objective of improving performance on downstream vision tasks. Our proposed GLOV employs a prompt search technique relying on embedding space guidance, that drives the prompt optimization for the VLMs. We iteratively query an LLM with downstream task-specific descriptions and ranked in-context examples derived from the previous (optimized) prompts. In-context examples guide the LLM toward the desired output, and their ranking (measured on a small held-out train set) provides the LLM with a prior for the language patterns preferred by the downstream VLM. To further steer the LLM generation towards a notion of goodness in each optimization step, we explicitly bias the language generation with a direction. This direction is determined by adding a hidden state offset vector (on the last token during the autoregressive generation) derived from the positive and negative prompts (based on their effectiveness on labeled training data) to the LLM’s activation space during generation. The intuition is that by directing the LLM generation toward the positive prompts, the model can discover semantically similar and potentially more effective solutions. One complete optimization run, depicting the effectiveness of the discovered solutions and the effect of applying the embedding space guidance is plotted in Figure 1. The best-performing prompts (on the held-out train set) achieve an absolute improvement of $2.6\%$ and $15.2\%$ on ImageNet (Deng et al., 2009) test set over CLIP (Radford et al., 2021) and LLaVA-OV (Li et al., 2024), respectively.
@@
We extensively evaluate our GLOV on the fundamental computer vision task of image classification and also on interpreting the safety and reliability of responses from open-source VLMs. For image classification, we demonstrate the generalization of our GLOV on a total of $16$ diverse datasets, with the two commonly employed families of VLMs – the dual-encoder and the recent visual encoder-decoder models (Radford et al., 2021; Li et al., 2024). On the other hand, for VLM safety, we evaluate our GLOV on $2$ recent benchmarks, which highlight severe vulnerabilities of open-source VLMs. Our GLOV can consistently discover highly effective solutions for the downstream task of interest resulting in significant improvements across the board. For example, the most effective prompts discovered for the dual-encoder models (e.g., CLIP) can improve the image classification accuracy by up to $15.0\%$ ($3.81\%$ on average) and for the encoder-decoder architectures (e.g., LLaVA), the resulting prompts show an even larger improvement of up to $57.5\%$ ($21.6\%$ on average). Furthermore, the effective prompts found for enhancing the VLM safety can lower the attack success rate (ASR) to merely $1.1\%$ (from $61.8\%$) with minimal degradation of the model’s generalization.
@@ 2 Related Work
Our work is related to large language and vision language models, approaches evaluating the safety of VLM responses, and prompt optimization methods (through LLMs) for VLMs.
@@ 2.1 LLMs and VLMs
Here, we first provide a brief overview of LLMs and then move towards VLMs.
@@
LLMs have revolutionized the natural language processing landscape. These models can typically be divided into two major groups; namely the long-short-term-memory (LSTM) (Hochreiter & Schmidhuber, 1997) and transformer-based architectures (Vaswani et al., 2017). The LSTM family of models uses gates to control the flow of information, allowing them to capture long-term dependencies in sequential data. Some notable works following the LSTM family of models include (Sutskever et al., 2014; Graves & Schmidhuber, 2005; Bahdanau et al., 2015; Beck et al., 2024). On the other hand, some notable works proposing the encoder-based transformer architectures include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019b), and DistilBERT (Sanh et al., 2020). The decoder-based LLMs, on the other hand, are designed for generative tasks such as text generation, translation, and summarization, with recent models including GPT-3 (Brown et al., 2020), T5 (Raffel et al., 2020), GPT-4 (OpenAI, 2023), and the Llama family of models (Dubey et al., 2024). In our work, since we need to access the weights of the LLMs, we resort to the open-source Llama-3 model. However, potentially any open-source LLM can be employed.
@@
VLMs can be placed in two categories. One group relies on dual-encoders (vision and text encoder), usually trained in a contrastive manner and these models are typically strong at tasks like image recognition. The most common among these methods are CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), OpenCLIP (Schuhmann et al., 2022), SigLIP (Zhai et al., 2023), and MetaCLIP (Xu et al., 2023). Many methods (Mirza et al., 2024, 2023b; Doveh et al., 2023b, a; Lin et al., 2023; Mirza et al., 2023a) build upon these models to further improve them for specific downstream tasks. The other group of methods aligns the visual modality with a frozen LLM and can be used for open-ended visual reasoning tasks like image captioning, visual question-answering, etc. Some representative approaches from this group include BLIP-2 (Li et al., 2023), Instruct-BLIP (Dai et al., 2023), MiniGPT (Zhu et al., 2024; Chen et al., 2024), and the LLaVA family of models (Liu et al., 2023; Li et al., 2024). Similarly, some approaches (Doveh et al., 2024; Gavrikov et al., 2024; Lin et al., 2024; Huang et al., 2024) build upon these models and provide further improvements. In our work, we focus on the task of object recognition by employing both families of models and frame the task of finding the optimal prompt templates (for CLIP (Radford et al., 2021)) and suitable prompts for open-ended generation (for LLaVA (Li et al., 2024)) as an optimization problem. Specifically, for the decoder-based VLMs, some recent works (Zhang et al., 2024) highlight that these models struggle for fine-grained object recognition. However, we show in our work, for the first time, that our GLOV can discover an optimal prompt that can greatly improve the visual recognition ability of these models, without requiring any gradient-based learning or fine-tuning.
2.2 Safety of VLMs
Enhancing the safety of open-source VLMs remains an open challenge. Recent works (Zong et al., 2024; Liu et al., 2024c; Tu et al., 2024; Gou et al., 2024) have evaluated multiple open-source VLMs on safety tasks. These tasks involve presenting the model with an unsafe image and an instruction, where the model is considered safe only if it refuses the unsafe instruction. However, current state-of-the-art VLMs exhibit significant vulnerabilities in this critical area. In our work, we enhance the safety of VLMs by optimizing for safety instructions and finding a general safety prompt that transcends models and safety benchmarks with minimal loss in the generalization abilities of the open-source VLMs. To the best of our knowledge, our GLOV is the first method to optimize safety instructions for VLMs, without any human in the loop.
@@ 2.3 Large-language Models as Prompt Optimizers
Some approaches propose employing LLMs (in an agentic workflow) to search for the optimal prompt for the downstream task. OPRO (Yang et al., 2024) coins the term “LLMs as Optimizers” and proposes iteratively discovering solutions (prompts) for natural language tasks by employing an LLM in a feedback loop. Similarly, Liu et al. (2024b) proposes to find suitable prompts for dual-encoder VLMs (e.g., CLIP) by iteratively prompting an LLM. Our GLOV also proposes to discover suitable prompts for VLMs but differs from Liu et al. (2024b) in the sense that we are employing a prompt that captures long-range dependencies by tapping into the history-buffer of the in-context examples and exploiting task-specific knowledge that helps to obtain prompts better suited for the downstream task. Furthermore, inspired by (Turner et al., 2023; Liu et al., 2024a) we propose a novel method for steering the LLM generation (through embedding space guidance) towards the text responses that are more suitable for downstream VLMs. Our GLOV– powered by a carefully tailored LLM prompt and the guidance scheme discovers highly effective solutions that help to enhance visual tasks for both the dual-encoder and encoder-decoder models.
3 GLOV
The goal of our GLOV is to improve the VLM’s downstream (vision) task performance by optimizing natural language prompts through employing an LLM in an iterative workflow. To achieve this, we build upon the successful line of work on in-context learning (Wei et al., 2022; Min et al., 2022), which can improve VLM performance through task-specific prompting (Mirza et al., 2024). However, our GLOV goes beyond these by automating the LLM-VLM interaction by adopting an agentic workflow and also explicitly guiding the language generation, conditioned on a prior of the difference of the sentence embeddings from the positive and negative prompts discovered during the previous optimization iterations. Although the application space of GLOV is general in terms of tasks and the downstream VLMs, for simplicity, here we focus our description around CLIP (Radford et al., 2021) while mentioning the differences for LLaVA (Li et al., 2024) and different tasks, where appropriate. Figure 2 provides an overview of our methodology.
For the ease of assimilation, we divide the description of our GLOV into different parts. In Section 3.1 we describe the fitness function and how it can provide an interface for the LLM-VLM interaction. In Section 3.2, we provide details about the tailored LLM prompt employed in our work. Finally, we conclude in Section 3.3 by providing details about the proposed guidance methodology.
@@ 3.1 LLM-VLM Interaction through Fitness Function
Figure 2: Overview of GLOV for dual-encoder models. GLOV consists of a specifically tailored LLM prompt, which constitutes system instruction, task description, and in-context examples (VLM prompts) which are evaluated (and ranked) on a few-shot training data in each iteration. This prompt instructs the LLM to generate several candidate solutions in each optimization iteration, conditioned on the in-context examples which are fed in conjunction with the accuracy values, highlighting their effectiveness. Furthermore, to steer the LLM generation towards the language preferred by the VLM, we add the scaled difference of the sentence embeddings (autoregressively) from the positive and negative text prompts to the intermediate layer of the LLM. This process is repeated until the stopping condition is met (e.g., maximum iterations). Note, that $H_{+}$ and $H_{-}$ refer to the sentence embeddings from the positive and negative text prompts.
The dataset-specific prompt templates $\mathcal{P}$ provided by CLIP (Radford et al., 2021) have been constructed manually, requiring human effort. In this work, we frame the prompt search as an optimization problem and propose to replace the human with an LLM, employed in an iterative feedback loop. Furthermore, we explicitly guide the generation process of the LLM in each optimization step by proposing a novel guidance methodology that can assist the LLM in understanding the style of language preferred by the downstream VLM, even though the two models only interact through a fitness function. At each optimization step $i$, the LLM provides multiple (e.g., $10$) solutions to improve the downstream task performance. However, not all solutions provided by the LLM are preferred (i.e., enhance performance) for the downstream vision task. To obtain a measure of the fitness (effectiveness) of the provided solutions to the downstream vision task, we evaluate all the candidate solutions on a held-out few-shot (1-shot) labeled training dataset $\mathcal{D}$. For CLIP (Radford et al., 2021), the zero-shot likelihood of class $\hat{c}$ for each discovered prompt $p\in\mathcal{P}$ during an optimization step can be found by
@@
$\text{Fitness}(p)=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}\left[\arg\max_{c}~{}~{}l_{c}(x)=y\right],$ (2)
where $\mathbbm{1}$ is an indicator function that is 1 if the predicted label matches the ground truth $y$ and 0 otherwise. For the encoder-decoder models (e.g., LLaVA), which produce open-ended outputs, we obtain the class likelihoods by obtaining a symbolic representation of the image in textual form and comparing the text embeddings from this symbolic representation with the text embeddings obtained for the individual class names, using a dedicated sentence embedding model (Reimers & Gurevych, 2019). We expand on these details in the Appendix Section A. Furthermore, for the open-ended task of enhancing the safety of VLMs, the fitness of the prompt is defined in terms of reducing the attack success rate (ASR) [^1] For safety enhancement, GLOV optimizes for a prompt that reduces the ASR on a downstream VLM.. To this end, we follow the evaluation protocols proposed by the original papers (Zong et al., 2024; Liu et al., 2024c), which focus on measuring the refusal rates and evaluations by using GPT-4 (OpenAI, 2023) as a judge for measuring the ASR.
It is important to note that the fitness function forms a bridge between two disjoint models – the LLM and the VLM, and is responsible for their interaction. The fitness (e.g., classification accuracy) provides feedback to the LLM regarding the type of natural language sentences that are preferred by the downstream VLM. The fitness function is responsible for ranking all the prompts provided as in-context examples to the LLM prompt in each optimization iteration (c.f., Section 3.2) and also forms the basis for the application of the embedding-space guidance methodology (c.f., Section 3.3) proposed in this work to bias the LLM responses towards a notion of goodness.
3.2 LLM Prompt
The LLM prompt (c.f., Appendix Figure 5) is responsible for driving the iterative prompt optimization. It consists of $3$ distinct parts, which are described as follows:
System prompt is a generic set of instructions that describe the task of the LLM. It helps a model to find the optimal prompts, improving the downstream task accuracy (or reducing ASR). The system prompt remains static for the entire optimization.